# MaLA-500: Massive Language Adaptation of Large Language Models

### v1

@@ Appendix B Detailed Results
ace_Latn 49.49 45.45 50.51 54.55 34.34 50.51 54.55 59.60 55.56 58.59 62.63 57.58 67.68 47.47
acm_Arab 45.45 51.52 65.66 64.65 34.34 50.51 53.54 52.53 55.56 56.57 57.58 57.58 55.56 76.77
afr_Latn 53.54 47.47 54.55 58.59 34.34 54.55 66.67 66.67 65.66 76.77 71.72 67.68 75.76 48.48
ajp_Arab 51.52 52.53 67.68 64.65 33.33 49.49 49.49 49.49 47.47 48.48 47.47 51.52 56.57 67.68
als_Latn 50.51 44.44 46.46 57.58 21.21 57.58 70.71 62.63 63.64 69.70 69.70 67.68 71.72 75.76
amh_Ethi 18.18 16.16 13.13 20.20 19.19 38.38 43.43 39.39 43.43 45.45 45.45 39.39 44.44 49.49
apc_Arab 46.46 55.56 69.70 64.65 25.25 50.51 48.48 47.47 45.45 55.56 51.52 52.53 53.54 55.56
arb_Arab 46.46 47.47 68.69 61.62 31.31 52.53 49.49 54.55 57.58 56.57 59.60 58.59 55.56 65.66
ary_Arab 42.42 42.42 67.68 61.62 24.24 46.46 47.47 45.45 44.44 49.49 45.45 46.46 50.51 74.75
arz_Arab 47.47 54.55 75.76 64.65 34.34 50.51 48.48 51.52 53.54 60.61 55.56 51.52 58.59 35.35
asm_Beng 32.32 47.47 63.64 54.55 22.22 41.41 48.48 47.47 47.47 47.47 49.49 46.46 52.53 64.65
ast_Latn 60.61 46.46 61.62 64.65 28.28 57.58 67.68 67.68 65.66 70.71 70.71 67.68 76.77 71.72
ayr_Latn 35.35 32.32 32.32 34.34 23.23 30.30 38.38 43.43 41.41 47.47 47.47 46.46 44.44 40.40
azb_Arab 40.40 37.37 37.37 41.41 27.27 38.38 38.38 39.39 45.45 51.52 40.40 43.43 45.45 70.71
azj_Latn 43.43 49.49 33.33 44.44 24.24 51.52 59.60 62.63 62.63 71.72 69.70 67.68 69.7 60.61
bak_Cyrl 35.35 54.55 34.34 44.44 34.34 49.49 61.62 63.64 63.64 70.71 64.65 63.64 65.66 72.73
bam_Latn 23.23 29.29 22.22 31.31 17.17 29.29 31.31 38.38 33.33 44.44 38.38 37.37 39.39 68.69
ban_Latn 57.58 52.53 56.57 60.61 25.25 43.43 59.60 62.63 59.60 61.62 63.64 59.6 61.62 64.65
bel_Cyrl 35.35 44.44 29.29 45.45 35.35 52.53 64.65 63.64 65.66 66.67 64.65 70.71 69.7 66.67
bem_Latn 29.29 31.31 35.35 38.38 23.23 33.33 42.42 44.44 48.48 46.46 44.44 49.49 47.47 55.56
ben_Beng 41.41 53.54 60.61 64.65 28.28 48.48 54.55 60.61 60.61 61.62 60.61 55.56 59.6 70.71
bjn_Latn 53.54 56.57 59.60 64.65 24.24 53.54 62.63 67.68 62.63 67.68 68.69 63.64 65.66 72.73
bod_Tibt 13.13 13.13 11.11 18.18 22.22 30.30 32.32 38.38 32.32 38.38 34.34 34.34 35.35 70.71
bos_Latn 65.66 54.55 46.46 55.56 27.27 59.60 69.70 64.65 64.65 70.71 75.76 68.69 75.76 72.73
bul_Cyrl 57.58 49.49 44.44 65.66 43.43 64.65 64.65 62.63 68.69 70.71 69.70 66.67 70.71 59.60
cat_Latn 61.62 52.53 57.58 66.67 29.29 57.58 68.69 61.62 65.66 68.69 69.70 66.67 72.73 66.67
ceb_Latn 61.62 48.48 48.48 49.49 29.29 61.62 69.70 73.74 66.67 80.81 74.75 72.73 74.75 60.61
ces_Latn 65.66 52.53 44.44 47.47 32.32 54.55 63.64 63.64 69.70 69.70 71.72 65.66 74.75 73.74
cjk_Latn 35.35 20.20 28.28 29.29 14.14 26.26 29.29 34.34 26.26 40.40 32.32 32.32 42.42 63.64
ckb_Arab 25.25 19.19 20.20 20.20 32.32 45.45 54.55 45.45 47.47 51.52 52.53 56.57 54.55 57.58
cmn_Hani 44.44 54.55 68.69 64.65 40.40 51.52 66.67 64.65 67.68 68.69 69.70 65.66 70.71 49.49
crh_Latn 42.42 43.43 42.42 50.51 25.25 55.56 53.54 59.60 52.53 63.64 57.58 52.53 58.59 77.78
cym_Latn 39.39 24.24 29.29 35.35 22.22 46.46 59.60 58.59 60.61 67.68 65.66 62.63 69.7 68.69
dan_Latn 60.61 53.54 53.54 61.62 31.31 52.53 65.66 65.66 67.68 74.75 74.75 71.72 73.74 34.34
deu_Latn 59.60 50.51 59.60 65.66 40.40 61.62 71.72 70.71 71.72 77.78 73.74 70.71 78.79 61.62
dyu_Latn 24.24 29.29 31.31 33.33 15.15 26.26 30.30 30.30 27.27 33.33 33.33 32.32 35.35 50.51
dzo_Tibt 16.16 12.12 15.15 17.17 19.19 32.32 28.28 28.28 28.28 29.29 31.31 35.35 32.32 72.73
ell_Grek 54.55 49.49 32.32 70.71 39.39 57.58 63.64 67.68 67.68 71.72 67.68 68.69 70.71 53.54
eng_Latn 66.67 48.48 66.67 71.72 44.44 68.69 72.73 71.72 73.74 75.76 76.77 76.77 77.78 76.77
epo_Latn 51.52 43.43 53.54 55.56 31.31 60.61 65.66 74.75 66.67 73.74 67.68 68.69 74.75 55.56
est_Latn 44.44 42.42 35.35 61.62 23.23 55.56 60.61 61.62 64.65 67.68 66.67 60.61 66.67 39.39
eus_Latn 45.45 46.46 58.59 61.62 22.22 49.49 45.45 57.58 54.55 58.59 56.57 53.54 57.58 59.60
ewe_Latn 23.23 26.26 25.25 38.38 21.21 36.36 37.37 38.38 31.31 37.37 42.42 39.39 37.37 56.57
fao_Latn 44.44 40.40 41.41 39.39 25.25 46.46 63.64 58.59 65.66 66.67 63.64 62.63 68.69 70.71
fij_Latn 30.30 30.30 33.33 42.42 17.17 34.34 41.41 45.45 46.46 50.51 46.46 45.45 48.48 39.39
fin_Latn 61.62 48.48 38.38 66.67 28.28 56.57 67.68 62.63 68.69 74.75 72.73 70.71 73.74 56.57
fon_Latn 22.22 25.25 26.26 30.30 20.20 31.31 32.32 28.28 30.30 34.34 31.31 34.34 34.34 75.76
fra_Latn 63.64 53.54 65.66 67.68 39.39 60.61 70.71 64.65 67.68 74.75 72.73 70.71 71.72 55.56
ful_Latn 35.35 31.31 34.34 44.44 18.18 34.34 44.44 41.41 45.45 49.49 56.57 46.46 58.59 59.60
fur_Latn 55.56 49.49 52.53 58.59 19.19 50.51 57.58 54.55 55.56 65.66 61.62 62.63 58.59 35.35
gla_Latn 30.30 27.27 23.23 34.34 15.15 31.31 40.40 36.36 34.34 36.36 39.39 39.39 43.43 69.70
gle_Latn 41.41 30.30 21.21 28.28 22.22 41.41 51.52 56.57 56.57 58.59 63.64 53.54 61.62 39.39
glg_Latn 56.57 45.45 63.64 65.66 31.31 54.55 69.70 64.65 72.73 75.76 78.79 74.75 81.82 41.41
grn_Latn 46.46 46.46 51.52 56.57 29.29 54.55 61.62 63.64 57.58 59.60 63.64 65.66 63.64 62.63
guj_Gujr 19.19 11.11 63.64 19.19 17.17 40.40 50.51 44.44 51.52 55.56 51.52 50.51 53.54 49.49
hat_Latn 45.45 37.37 42.42 64.65 27.27 52.53 58.59 62.63 63.64 64.65 70.71 60.61 75.76 45.45
hau_Latn 26.26 29.29 27.27 38.38 19.19 49.49 54.55 55.56 54.55 54.55 55.56 53.54 55.56 40.40
heb_Hebr 42.42 44.44 21.21 26.26 20.20 36.36 35.35 40.40 35.35 31.31 40.40 40.4 35.35 63.64
hin_Deva 51.52 51.52 63.64 70.71 24.24 42.42 53.54 58.59 57.58 61.62 60.61 57.58 56.57 65.66
hne_Deva 44.44 48.48 66.67 64.65 28.28 49.49 52.53 60.61 59.60 59.60 59.60 59.6 55.56 56.57
hrv_Latn 66.67 56.57 52.53 51.52 30.30 56.57 73.74 67.68 71.72 69.70 76.77 69.7 76.77 74.75
hun_Latn 65.66 51.52 36.36 43.43 33.33 45.45 57.58 59.60 56.57 69.70 61.62 59.6 65.66 63.64
hye_Armn 29.29 13.13 18.18 17.17 26.26 52.53 59.60 60.61 67.68 70.71 66.67 67.68 71.72 71.72
ibo_Latn 20.20 27.27 42.42 28.28 31.31 50.51 57.58 55.56 48.48 61.62 58.59 51.52 53.54 66.67
ilo_Latn 47.47 49.49 51.52 53.54 19.19 51.52 62.63 65.66 60.61 68.69 72.73 68.69 67.68 73.74
ind_Latn 63.64 52.53 62.63 71.72 27.27 56.57 72.73 69.70 69.70 77.78 73.74 76.77 77.78 57.58
isl_Latn 48.48 37.37 36.36 34.34 25.25 46.46 56.57 53.54 59.60 54.55 66.67 66.67 67.68 55.56
ita_Latn 60.61 54.55 61.62 64.65 35.35 59.60 71.72 66.67 67.68 72.73 76.77 71.72 73.74 70.71
jav_Latn 51.52 58.59 60.61 61.62 31.31 46.46 63.64 63.64 64.65 64.65 69.70 64.65 67.68 73.74
jpn_Jpan 41.41 51.52 64.65 65.66 36.36 56.57 64.65 67.68 66.67 75.76 74.75 69.7 72.73 56.57
kab_Latn 14.14 19.19 17.17 27.27 14.14 31.31 32.32 21.21 26.26 30.30 26.26 25.25 29.29 66.67
kac_Latn 32.32 26.26 34.34 35.35 21.21 34.34 35.35 40.40 39.39 42.42 39.39 41.41 39.39 63.64
kam_Latn 29.29 32.32 32.32 44.44 20.20 29.29 42.42 42.42 39.39 49.49 36.36 40.4 42.42 41.41
kan_Knda 15.15 10.10 59.60 19.19 23.23 45.45 48.48 55.56 56.57 58.59 54.55 55.56 56.57 39.39
kat_Geor 42.42 50.51 20.20 26.26 24.24 45.45 56.57 61.62 59.60 61.62 67.68 62.63 66.67 44.44
kaz_Cyrl 36.36 52.53 35.35 49.49 27.27 47.47 59.60 62.63 62.63 67.68 66.67 64.65 63.64 58.59
kbp_Latn 22.22 25.25 27.27 26.26 21.21 30.30 30.30 38.38 31.31 39.39 37.37 42.42 41.41 51.52
kea_Latn 51.52 39.39 57.58 57.58 29.29 52.53 58.59 60.61 57.58 65.66 64.65 66.67 70.71 64.65
khm_Khmr 31.31 13.13 30.30 29.29 38.38 53.54 58.59 55.56 58.59 65.66 60.61 63.64 67.68 67.68
kik_Latn 28.28 33.33 34.34 31.31 21.21 41.41 39.39 46.46 43.43 50.51 46.46 45.45 50.51 47.47
kin_Latn 31.31 32.32 49.49 30.30 18.18 32.32 44.44 49.49 50.51 52.53 49.49 48.48 54.55 36.36
kir_Cyrl 38.38 50.51 32.32 47.47 29.29 52.53 59.60 61.62 63.64 69.70 60.61 62.63 66.67 71.72
kmb_Latn 28.28 34.34 34.34 43.43 15.15 28.28 41.41 39.39 37.37 39.39 39.39 44.44 42.42 52.53
kmr_Latn 30.30 34.34 38.38 42.42 20.20 45.45 52.53 59.60 57.58 58.59 54.55 54.55 62.63 69.70
kon_Latn 34.34 35.35 44.44 46.46 26.26 43.43 46.46 56.57 51.52 50.51 55.56 48.48 53.54 70.71
kor_Hang 48.48 53.54 44.44 60.61 42.42 62.63 67.68 69.70 68.69 76.77 77.78 69.7 74.75 43.43
lao_Laoo 25.25 10.10 20.20 31.31 31.31 49.49 56.57 61.62 60.61 66.67 63.64 59.6 62.63 58.59
lij_Latn 52.53 51.52 51.52 60.61 23.23 51.52 66.67 64.65 63.64 64.65 68.69 68.69 69.7 54.55
Table 5: Detailed Results (Part I). For previous LLMs, 3-shot results are presented.

### v2

@@ Appendix B Detailed Results
ace_Latn 44.12 47.55 50.00 36.76 34.31 52.94 60.29 60.78 65.69 67.65 64.22 65.20 68.63 71.57
acm_Arab 52.45 65.69 69.12 58.33 32.35 53.43 59.31 63.73 63.73 67.16 66.67 69.12 66.67 66.67
afr_Latn 68.14 55.39 53.92 40.20 41.18 62.25 65.69 69.12 71.08 74.02 73.53 74.51 76.96 78.92
ajp_Arab 47.55 64.22 68.63 53.43 33.33 56.86 59.80 59.80 65.20 63.73 63.24 69.12 68.14 66.67
als_Latn 41.67 46.57 45.59 28.43 27.94 51.96 63.73 62.25 69.12 71.08 69.61 72.06 75.98 77.45
amh_Ethi 15.69 18.63 16.67 13.24 25.00 36.76 45.59 51.47 51.96 53.92 51.96 53.43 54.90 53.92
apc_Arab 46.57 65.69 68.14 53.43 31.37 55.88 60.29 65.69 65.69 67.16 65.20 68.63 67.65 72.06
arb_Arab 53.43 63.24 68.14 57.35 32.35 54.90 60.29 63.73 65.20 68.14 67.16 69.12 68.63 70.59
ary_Arab 45.10 57.84 69.12 50.49 26.47 52.45 55.39 56.37 60.29 59.80 64.22 63.73 59.31 64.71
arz_Arab 50.98 64.22 68.14 56.86 30.88 52.45 59.31 60.78 64.22 66.18 68.14 69.12 66.67 69.12
asm_Beng 17.16 49.02 61.27 37.25 31.37 53.43 58.82 65.20 67.65 67.65 68.14 67.65 67.65 67.65
ast_Latn 69.12 60.78 69.12 55.39 34.31 65.69 70.10 70.59 74.02 75.00 75.98 77.94 79.90 79.90
ayr_Latn 25.00 26.96 32.35 19.61 20.10 29.41 38.24 38.73 38.24 43.14 40.20 43.14 44.61 42.16
azb_Arab 25.49 41.67 32.84 24.51 25.98 41.18 45.59 45.10 46.57 49.51 50.00 49.02 50.49 49.51
azj_Latn 34.80 64.22 37.25 32.84 30.88 57.84 64.71 68.63 64.22 72.55 69.61 70.59 74.02 72.55
bak_Cyrl 38.73 61.27 32.35 32.35 34.80 51.47 60.29 61.27 69.12 68.63 68.14 68.63 73.53 70.10
bam_Latn 25.49 24.51 29.41 20.10 22.55 25.98 34.80 42.16 43.14 44.12 45.10 42.16 46.08 44.12
ban_Latn 58.82 51.47 58.82 43.14 28.92 55.39 63.24 65.69 66.67 72.06 72.06 72.55 72.06 71.57
bel_Cyrl 47.55 59.80 28.92 30.39 40.69 60.29 63.24 66.18 67.65 70.10 72.55 72.06 72.55 73.04
bem_Latn 31.37 28.92 38.24 25.49 21.08 34.80 43.14 48.04 50.49 50.49 53.43 53.43 52.45 53.92
ben_Beng 25.49 61.27 64.22 52.45 31.37 54.90 63.24 62.25 67.65 70.10 70.10 69.12 66.18 68.63
bjn_Latn 48.53 51.96 61.76 42.65 32.35 62.75 66.18 68.14 71.57 75.98 73.04 72.55 75.00 77.45
bod_Tibt 15.20 12.75 15.69 15.69 22.06 34.80 37.75 37.75 38.73 39.71 39.71 39.71 41.67 44.12
bos_Latn 65.20 64.71 45.59 33.82 37.75 65.20 72.06 70.10 71.57 75.98 75.00 76.47 76.96 77.45
bul_Cyrl 66.18 63.24 38.73 52.94 45.10 62.25 67.65 67.16 68.14 75.00 71.57 72.06 73.53 75.00
cat_Latn 71.08 60.78 66.67 60.29 34.31 59.31 68.14 68.63 71.57 72.55 69.12 73.04 76.96 76.47
ceb_Latn 50.49 50.98 49.02 39.71 39.22 60.29 66.67 66.67 68.63 73.04 72.06 71.57 74.51 74.02
ces_Latn 69.12 62.75 47.55 40.69 39.22 62.25 69.61 70.59 72.55 76.47 72.55 74.02 80.88 76.96
cjk_Latn 27.94 30.39 34.31 26.47 22.55 30.88 31.86 32.84 38.24 38.24 38.24 35.78 39.22 42.65
ckb_Arab 19.61 22.55 23.04 12.25 28.92 53.43 60.29 57.35 65.20 65.20 62.75 65.69 65.69 70.59
cmn_Hani 73.04 65.20 67.65 54.90 39.71 69.12 74.02 72.06 76.47 77.45 76.47 75.98 79.41 76.47
crh_Latn 38.24 56.37 40.20 36.76 29.41 51.96 62.25 61.76 64.22 69.12 64.22 70.10 69.12 71.08
cym_Latn 39.22 28.43 34.80 21.57 28.43 55.39 62.75 63.73 66.18 72.06 74.02 72.06 75.00 77.45
dan_Latn 69.12 64.22 55.39 44.12 38.24 54.41 63.24 65.20 70.10 71.08 72.06 71.57 73.53 74.51
deu_Latn 74.02 60.29 61.27 55.39 41.18 63.73 68.63 71.57 69.12 75.00 75.49 76.47 77.45 77.45
dyu_Latn 28.43 28.92 32.35 20.10 21.08 29.90 38.73 39.71 46.57 44.12 41.67 47.06 44.61 43.63
dzo_Tibt 14.71 10.29 13.73 12.75 21.57 30.39 36.76 37.25 39.71 36.76 39.22 37.75 43.14 39.22
ell_Grek 47.55 63.73 28.43 60.29 43.63 62.75 69.61 66.67 69.12 69.61 70.59 73.04 72.06 72.06
eng_Latn 71.57 59.80 71.08 67.65 48.04 63.24 70.59 69.12 69.12 74.02 73.04 74.51 76.96 75.98
epo_Latn 52.94 50.49 52.94 43.63 27.94 49.51 66.18 66.67 68.63 72.55 73.53 71.57 74.51 75.98
est_Latn 48.04 54.90 41.18 57.35 29.41 55.88 62.75 66.67 70.10 72.06 70.10 69.12 71.57 73.04
eus_Latn 36.27 59.80 64.22 55.88 27.94 52.45 61.76 66.18 66.18 73.53 73.53 72.55 73.53 75.49
ewe_Latn 23.53 23.53 29.90 17.16 23.53 28.43 38.24 35.29 41.18 43.63 38.73 44.61 39.71 43.63
fao_Latn 41.18 43.63 38.73 29.90 35.29 52.94 56.86 57.84 62.25 61.27 62.25 61.76 64.22 69.12
fij_Latn 27.45 27.45 36.27 24.02 24.51 37.75 48.53 47.55 53.92 48.53 50.00 51.47 51.47 53.92
fin_Latn 67.65 63.24 38.73 56.86 36.76 61.76 70.10 70.10 72.55 74.51 72.06 73.53 75.00 75.49
fon_Latn 25.49 22.06 31.37 19.12 23.53 29.90 30.88 37.25 35.78 38.24 39.71 38.24 37.25 46.57
fra_Latn 72.06 64.71 66.18 59.80 36.76 58.82 71.08 67.65 71.08 74.51 71.57 74.02 77.45 77.45
ful_Latn 27.45 31.37 32.84 24.02 21.57 31.86 36.76 40.69 43.14 45.10 44.12 47.06 47.06 46.08
fur_Latn 58.82 50.00 55.88 38.73 35.78 53.92 55.88 62.25 66.67 68.63 68.14 67.16 75.00 72.55
gla_Latn 37.75 24.51 27.45 17.16 24.51 50.98 55.88 57.84 57.35 59.80 63.24 61.27 62.75 65.20
gle_Latn 39.71 25.49 25.00 17.16 27.94 53.43 57.84 63.24 64.22 67.65 64.71 63.24 62.75 73.53
glg_Latn 68.63 62.25 64.22 55.39 29.41 66.67 70.10 71.08 74.02 76.96 73.53 76.47 77.45 80.39
grn_Latn 42.16 47.06 48.53 33.82 25.98 52.45 62.75 64.71 62.25 67.16 64.22 65.69 61.76 69.61
guj_Gujr 15.20 09.31 62.25 12.75 28.43 50.98 54.90 60.29 63.73 63.73 64.22 65.69 62.75 67.16
hat_Latn 41.67 38.73 42.16 45.10 34.80 57.35 63.73 62.25 65.20 72.06 69.61 70.10 73.53 73.04
hau_Latn 25.49 28.43 29.90 20.59 28.43 47.06 55.39 57.84 61.27 65.69 62.75 65.69 66.18 65.20
heb_Hebr 37.75 63.24 20.59 11.76 26.47 41.67 47.06 51.47 54.90 51.96 50.98 54.90 53.92 54.41
hin_Deva 44.61 62.75 62.75 51.96 33.82 55.39 60.78 66.67 65.20 69.61 70.59 74.02 73.04 72.06
hne_Deva 37.75 58.82 59.80 49.02 28.43 55.39 55.88 65.20 62.75 68.63 66.18 65.69 68.14 71.08
hrv_Latn 66.18 65.20 44.12 36.27 40.69 63.73 73.04 71.08 73.53 76.47 72.06 74.51 78.43 78.43
hun_Latn 71.08 63.24 41.67 27.94 30.88 60.78 67.16 70.59 68.63 75.49 73.53 74.02 73.04 76.47
hye_Armn 20.59 17.16 13.73 12.75 32.84 58.82 59.80 67.16 65.20 69.61 68.14 69.12 69.12 72.55
ibo_Latn 24.02 26.47 38.24 19.12 30.39 51.47 57.35 63.73 67.16 69.12 68.14 69.12 68.63 72.06
ilo_Latn 45.10 45.59 48.04 32.35 27.45 54.90 61.76 61.76 68.14 68.14 70.10 73.04 73.53 70.10
ind_Latn 74.02 62.75 70.10 54.90 40.69 62.75 68.63 71.57 70.59 75.49 75.49 76.96 80.39 77.94
isl_Latn 35.29 36.76 28.92 24.51 38.73 55.88 60.78 58.33 60.29 64.71 63.73 63.24 62.75 65.20
ita_Latn 69.61 62.25 62.75 57.84 40.20 64.22 70.59 70.59 74.51 77.94 75.98 76.96 80.39 76.96
jav_Latn 50.49 52.94 55.39 38.24 31.86 53.43 60.78 64.22 65.20 72.55 69.12 73.04 68.63 73.04
jpn_Jpan 73.53 60.29 63.24 55.88 38.73 67.16 72.06 75.49 78.92 79.41 80.39 78.92 81.86 81.37
kab_Latn 16.18 16.67 20.10 12.25 20.59 24.02 22.55 30.39 31.86 34.80 28.43 33.33 32.35 34.31
kac_Latn 25.98 24.51 28.43 20.59 20.10 24.02 29.90 35.78 35.78 43.14 37.75 37.25 43.63 39.71
kam_Latn 26.96 34.31 34.80 26.47 22.06 36.76 38.73 37.75 40.69 41.67 46.57 41.67 42.16 42.16
kan_Knda 17.16 11.27 61.27 11.27 25.49 50.98 57.35 60.29 61.27 65.20 63.24 64.22 65.69 67.16
kat_Geor 29.41 61.27 18.14 14.71 32.84 56.86 60.78 62.25 65.20 67.65 70.59 70.59 70.10 74.51
kaz_Cyrl 37.75 62.75 29.90 28.43 34.31 53.43 57.35 62.25 65.69 67.16 65.20 65.69 69.12 67.65
kbp_Latn 24.51 22.06 30.39 16.18 21.08 28.43 36.76 38.73 40.69 39.22 40.69 39.71 38.73 40.20
kea_Latn 53.43 51.96 56.86 39.71 32.84 56.86 63.73 65.20 67.16 69.61 69.12 71.57 71.57 72.06
khm_Khmr 27.45 11.27 25.49 15.20 39.22 61.76 67.16 67.65 68.63 72.06 73.53 75.00 76.47 76.47
kik_Latn 29.41 32.84 38.73 26.96 21.57 37.75 49.51 50.49 50.49 56.86 56.37 56.37 52.94 56.86
kin_Latn 26.47 32.35 50.49 24.51 27.45 40.69 49.02 52.94 57.84 58.33 60.78 56.86 59.31 59.80
kir_Cyrl 35.78 60.78 34.80 27.45 29.90 45.59 58.33 60.78 60.29 65.20 60.29 64.71 66.18 66.18
kmb_Latn 26.47 28.43 33.82 25.00 21.08 31.86 35.29 39.71 38.24 41.18 41.67 37.25 41.67 44.61
kmr_Latn 29.41 33.82 33.33 21.57 25.98 37.75 47.06 47.55 52.45 52.45 54.90 54.41 58.82 61.76
kon_Latn 33.33 33.82 40.69 32.35 22.06 39.71 46.57 51.96 53.92 64.71 64.22 60.78 64.22 65.20
kor_Hang 67.65 63.24 43.14 56.37 45.10 63.24 67.65 69.12 71.57 73.04 70.59 75.98 76.96 76.47
lao_Laoo 24.02 14.22 26.47 16.67 39.71 55.39 62.25 63.73 68.63 70.59 70.10 68.63 70.59 70.59
lij_Latn 55.88 53.43 56.37 44.61 37.25 58.82 67.65 66.67 69.61 71.57 71.08 74.02 76.47 74.02
Table 22: Detailed results on SIB200 (Part I). For previous LLMs, 3-shot results are presented.